Hi can someone please help
Find the measure of an angle whose complement is twenty nine times its measure.

Answer by Theo(13342)

the angles are complements so their sum is 90 degrees.

let the first angle be x.
let the second angle be y.

you get x + y = 90

since y = 29 * x, we can replace y with 29 * x and the formula becomes x + 29x = 90.

combine like terms to get 30x = 90

divide both sides of the equation by 30 to get x = 3

y is equal to 29 * x, so y = 29 * 3 = 87.

x + y becomes 3 + 87 which becomes 90.

x and y are complements because their sum is 90.

your solution is that the angle whose complement is 29 times its measure is 3 degrees.